Moolah has 27 coins that are all dimes and quarters. The value of the coins is $4.35. How many dimes and how many quarters does Moolah have?
Answer by checkley77(12569)

D+Q=27 MULTIPLY BY -.10 & ADD.
.10D+.25Q=4.35
-.10D-.10Q=-2.7
---------------------
.15Q=1.65
Q=1.65/.15
Q=11 QUARTERS.
D+11=27
D=27-11
D=16 DIMES.
PROOF:
.10*16+.25*11=4.35
1.60+2.75=4.35
4.35=4.35